Most communities have a Sign Ordinance that limits the amount of signage that you use to advertise your business. Often this is computed by looking at factors such as highway frontage and size of your building.
Suppose that your sign is limited to 3600 square inches (width x height = 3600) . You want to build a sign that is aesthetically appealing to advertise your business. To be an aesthetically appealing sign you would want the width to be 2 times the height and you want the top and bottom margins to both be 5 inches each and the right and left margin to both be 7 inches each. The "Printable Region" would be the area inside these margins and this will be the area that you can use to advertise your business.


How many square inches will be in the "Printable Region"

Answer by Boreal(15235)

height=h
width=2h
the printable area is 2h-14 (width) and the height is h-10 inches
The product h*2h=2h^2=3600
h^2=1800
h=30 sqrt(2)
Therefore, the printable area is 60(sqrt(2))-14 and 30 sqrt(2)-10
That product is 1800*2-1020* sqrt(2)+140=3740-1020 sqrt(2) sq. inches or numerically 2297.5 sq inches
Check is the sign is roughly 42.4 X 84.8
The height shrinks to 32.4 and the width shrinks to 70.8. That product is 2293.9 or in the range.